# Dynamical phases of short-term memory mechanisms in RNNs

### v1

@@ Acknowledgements
We would like to thank members of the Geometric Intelligence Lab for their helpful feedback on an earlier version of the project and Dr. Boris Shraiman for fruitful discussions on the final manuscript. EC and BK’s internships were supported in part by a grant from the Feldman-McClelland Open-a-Door fund of the Pittsburgh Foundation. BK thanks the Impact Scholarship and Research Scholarship for Critical Thinkers programs from the Bridge to Turkiye Funds for supporting his visit to Stanford University. FD expresses gratitude for the valuable mentorship he received at PHI Lab during his internship at NTT Research. MJS gratefully acknowledges funding from the Simons Collaboration on the Global Brain and the Vannevar Bush Faculty Fellowship Program of the U.S. Department of Defense. NM acknowledges funding from the National Science Foundation, award 2313150. FD acknowledges funding from Stanford University’s Mind, Brain, Computation and Technology program, which is supported by the Stanford Wu Tsai Neuroscience Institute. This research was supported in part by grant NSF PHY-2309135 and the Gordon and Betty Moore Foundation Grant No. 2919.02 to the Kavli Institute for Theoretical Physics (KITP). Some of the computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results.
Appendix S1 Details on the toy models
@@
S1.1 The ghost model
The first model we considered is the ghost model, which effectively forms a slow point that can be used to store the duration of the delay before the network takes an action. This model was introduced in a prior work [37], whose results we briefly summarize below. Then, we generalize the methodology introduced in the earlier work, which we will eventually use to study the additional toy models we introduced in this work.
S1.1.1 Summary of the previous work
@@
$\tau\dot{x}(t)=x(t)^{2}+r,$ (S1)
where $x\in\mathbb{R}$ is the state variable and $r\in\mathbb{R}$ is a learnable parameter. Here, $\tau$ is a fixed time scale, which we effectively set to one such that time is computed in dimensionless units (in units of $\tau$). This system is very well known in the traditional dynamical system theory literature as the standard form for the saddle-node bifurcation [41]. Specifically, for $r<0$, the dynamical system has two fixed points $x^{*}=\pm\sqrt{-r}$, the one on the left being attractive and the other repulsive. However, for $r>0$, the system has no fixed-points. Hence, a “saddle-node bifurcation” is said to take place as $r$ is varied between positive and negative values [41]. Since local minima of most functions can be approximated by a quadratic term, this is considered the standard form for the saddle-node bifurcation and can be used to approximate the emergence/collusion of two fixed-points.
Though traditional work often focuses on changes as $r$ is varied by hand, [37] has taken an alternative approach and considered learning this parameter through a task. Specifically, the output of the network is defined as $\hat{o}(x)=\Theta(x-x^{*})$, where $\Theta(\cdot)$ is the Heaviside function and $x^{*}$ is some fixed-parameter (which is later taken to $\infty$ in analytical calculations). The model is initialized at $x(t=0)=0$. Then, the goal of the model is to output 00 until some time $T$ and then $1$ for another $T$ times. In mathematical terms, $o(t)=\Theta(t-T)$. Unlike the prior work [37], since our goal is to understand the scaling of the effective learning rates with respect to the delay period lengths, we define a normalized loss function as:
@@
It is possible to compute this loss function exactly in the limit $x^{*}\to\infty$, which is as follows [37]:
@@
These expressions differ from [37] only in the sense that the normalization of the loss function introduces a $(T)^{-1}$ coefficient to the gradient computation. Notably, this loss function has a kink at its global optimum. Thanks to this kink, [37] has defined a maximum (effective) learning rate $\alpha^{*}$, for any $\alpha>\alpha^{*}$, the model would be thrown from the global minimum to the zero gradient regime, after which the learning would come to an halt despite a non-desirable loss value of $1$. This learning rate can be computed as:
@@
S1.1.2 Ghost model with asymmetric response time
@@
As shown by previous work [37] (by integrating the dynamical system equations), in the limit $x^{*}\to\infty$, the network output can be written as $\hat{o}(t)=\Theta(t-t^{*})$, where $t^{*}=\frac{\pi}{2\sqrt{r}}$. This creates three distinct regimes of analytical calculation: i) $t^{*}\leq T_{\rm delay}$, ii) $T_{\rm delay}\leq t^{*}\leq T_{\rm delay}+T_{\rm resp}$, and iii) $T_{\rm delay}+T_{\rm resp}\leq t^{*}$.
@@
$\mathcal{L}(r)=\begin{cases}1\quad&\text{for}\quad r\leq\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm res})^{2}},\\ \frac{\pi}{2T_{\rm resp}\sqrt{r}}-\frac{T_{\rm delay}}{T_{\rm resp}}\quad&\text{for}\quad\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm resp})^{2}}\leq r\leq\frac{\pi^{2}}{4T_{\rm delay}^{2}},\\ 1-\frac{\pi}{2T_{\rm delay}\sqrt{r}}\quad&\text{for}\quad r\geq\frac{\pi^{2}}{4T_{\rm delay}^{2}}.\end{cases}$ (S11)
@@
Similar to before, one can compute the critical learning rate due to the kink at the global optimum. First, noting $r_{b}=\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm res})^{2}}$, we arrive at:
$r^{*}-r_{b}=\frac{\pi^{2}}{4T_{\rm delay}^{2}}-\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm res})^{2}}=\frac{\pi^{2}(2T_{\rm delay}+T_{\rm resp})T_{\rm resp}}{4T_{\rm delay}^{2}(T_{\rm delay}+T_{\rm res})^{2}}\xrightarrow{T_{\rm delay}\gg T_{\rm delay}}\frac{\pi^{2}T_{\rm resp}}{2T_{\rm delay}^{3}}.$ (S13)
@@
S1.2 Limit cycle model
@@
S1.2.2 Analytical calculations with equal delay and response times
@@
S1.2.3 Analytical calculations with long delays
@@
Appendix S2 Methods
S2.1 Metrics
We designed two metrics—reaction accuracy, reaction reliability—to evaluate the performance of our RNN models. While reaction accuracy considers only the maximum output channel, reaction reliability takes into account the model’s confidence in its output. To explain this more formally, let $V$ be the output vector in the reaction interval, where $V\in\mathbb{R}^{N\times 2}$, and let $G$ be the ground truth vector in the reaction interval, where $G\in\mathbb{R}^{N\times 1}$.
Reaction Accuracy
$RA=1-\left|\arg\max(V)-G\right|$ (S30)
Reaction Reliability
$RR=\frac{1}{N}\sum_{i=1}^{N}V_{i,G_{i}}$ (S31)
Reaction Accuracy (RA) is defined using the equation where $\arg\max(V)$ gives the index of the maximum value in $V$ for each instance. Reaction Reliability (RR) is calculated as the mean of the probabilities corresponding to the ground truth values $G$ for each instance.
S2.2 Algorithm Detection in Task-Trained RNNs
To manage the large-scale results from RNN evaluations, we developed a rule-based algorithm classification method. This subsection explains the criteria used for classification and the corresponding subsets.
Oscillation Measure
$\mathcal{OM}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(V_{i}-\bar{V}\right)^{2}}$ (S32)
where $V$ represents the firing rates in the post-reaction period’s last $N$ time steps, with $N$ corresponding to the base trial length, i.e., the sum of the input interval, delay interval, and reaction interval lengths ($V\in\mathbb{R}^{N}$), and $\bar{V}$ is the mean of these values.

### v2

@@ Acknowledgements
We would like to thank members of the Geometric Intelligence Lab for their helpful feedback on an earlier version of the project and Dr. Boris Shraiman for fruitful discussions on the final manuscript. EC and BK’s internships were supported in part by a grant from the Feldman-McClelland Open-a-Door fund of the Pittsburgh Foundation. BK thanks the Impact Scholarship and Research Scholarship for Critical Thinkers programs from the Bridge to Turkiye Funds for supporting his visit to Stanford University. FD expresses gratitude for the valuable mentorship he received at PHI Lab during his internship at NTT Research and acknowledges funding from Stanford University’s Mind, Brain, Computation and Technology program, which is supported by the Stanford Wu Tsai Neuroscience Institute.. MJS gratefully acknowledges funding from the Simons Collaboration on the Global Brain and the Vannevar Bush Faculty Fellowship Program of the U.S. Department of Defense. NM acknowledges funding from the National Science Foundation, award 2313150. PY was additionally supported by the Foundation of the Shanghai Municipal Education Commission (No. 24RGZNA01). This research was supported in part by grant NSF PHY-2309135 and the Gordon and Betty Moore Foundation Grant No. 2919.02 to the Kavli Institute for Theoretical Physics (KITP). Some of the computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results.
Impact Statement
This paper contributes to machine learning and computational neuroscience by making an unprecedented number of trained RNNs publicly available—models that required several weeks of compute time to generate.
Appendix S1 Theoretical details on attractor mechanisms subserving short-term memory
S1.1 A short background on dynamical system theory
A dynamical system describes how the state of a system evolves over time. Formally, it can be represented by a differential equation of the form:
$\dot{x}(t)=f(x(t)),\quad x(t_{0})=x_{0},$
where $x_{0}$ denotes the initial state, and $\dot{x}$ represents the time derivative of the state variable $x\in\mathbb{R}^{n}$.
In neuroscience, researchers often examine neural firing rates $r\in\mathbb{R}^{n}$ through the lens of dynamical systems. This approach helps uncover how temporal dynamics unfold within lower-dimensional neural state spaces, as characterized by dynamical systems theory.
Since firing rates are themselves components of the system’s state, in this study we simulate such systems using leaky firing-rate RNNs—biologically interpretable neural networks. One might argue that RNNs may not be expressive enough to learn such complex dynamics and that more sophisticated architectures like Long Short-Term Memory networks (LSTMs) or Transformers might perform better. However, this concern is addressed by the universal approximation property of RNNs, which states that, given a sufficient number of neurons, even vanilla RNNs can approximate any dynamical system (Schäfer & Zimmermann, 2006). Additionally, compared to more complex models, RNNs offer greater interpretability due to their simpler and shallower architecture.
As previously mentioned, neuroscientists often study neural activity through the lens of attractor dynamics. In this framework, researchers analyze how systems evolve over time by characterizing their underlying attractor structures. As we focus on several commonly studied attractors from dynamical systems theory (Strogatz, 2018), we describe the fixed-point attractors, slow-point manifolds, and limit cycles below.
Fixed-Point Attractors:
In a fixed-point attractor, the system evolves toward a specific state where it remains indefinitely. Once the state is reached, the time derivative becomes zero, i.e., $\dot{x}=0$.
Slow-Point Manifolds:
Slow-point manifolds are particularly relevant to memory-related mechanisms in neuroscience. These manifolds describe regions of the state space where the system changes very slowly, satisfying $\dot{x}\approx 0$. Unlike fixed points, which trap the system, slow points allow it to persist along a trajectory over time. This property enables the system to maintain information temporarily, as illustrated in Fig. S4D.
Limit Cycle Attractors:
Limit cycle attractors describe systems that exhibit sustained oscillatory behavior. These trajectories form closed loops in state space and are often associated with rhythmic or repeated tasks. In such cases, $\dot{x}\neq 0$, as shown in Fig. S4C. However, as demonstrated in our experiments, limit cycles can closely approximate slow-point manifolds during the trial period.
Importantly, a single dynamical system may contain multiple attractor types, and the specific trajectory it follows can depend on the initial condition $x_{0}$. Nonetheless, identifying and understanding these fundamental attractor structures is essential for uncovering the mechanisms underlying neural population dynamics.
Appendix S2 Theoretical studies of scaling relationships via toy models
@@
S2.1 The ghost model
The first model we considered is the ghost model, which effectively forms a slow point that can be used to store the duration of the delay before the network takes an action. This model was introduced in a prior work (Dinc et al., 2025b), whose results we briefly summarize below. Then, we generalize the methodology introduced in the earlier work, which we will eventually use to study the additional toy models we introduced in this work.
S2.1.1 Summary of the previous work
@@
$\tau\dot{x}(t)=x^{2}(t)+r,$ (S1)
where $x\in\mathbb{R}$ is the state variable and $r\in\mathbb{R}$ is a learnable parameter. Here, $\tau$ is a fixed time scale, which we effectively set to one such that time is computed in dimensionless units (in units of $\tau$). This system is very well known in the traditional dynamical system theory literature as the standard form for the saddle-node bifurcation (Strogatz, 2018). Specifically, for $r<0$, the dynamical system has two fixed points $x^{*}=\pm\sqrt{-r}$, the one on the left being attractive and the other repulsive. However, for $r>0$, the system has no fixed-points. Hence, a “saddle-node bifurcation” is said to take place as $r$ is varied between positive and negative values (Strogatz, 2018). Since local minima of most functions can be approximated by a quadratic term, this is considered the standard form for the saddle-node bifurcation and can be used to approximate the emergence/collusion of two fixed-points.
Though traditional work often focuses on changes as $r$ is varied by hand, (Dinc et al., 2025b) has taken an alternative approach and considered learning this parameter through a task. Specifically, the output of the network is defined as $\hat{o}(x)=\Theta(x-x^{*})$, where $\Theta(\cdot)$ is the Heaviside function and $x^{*}$ is some fixed-parameter (which is later taken to $\infty$ in analytical calculations). The model is initialized at $x(t=0)=0$. Then, the goal of the model is to output 00 until some time $T$ and then $1$ for another $T$ times. In mathematical terms, $o(t)=\Theta(t-T)$. Unlike the prior work (Dinc et al., 2025b), since our goal is to understand the scaling of the effective learning rates with respect to the delay period lengths, we define a normalized loss function as:
@@
It is possible to compute this loss function exactly in the limit $x^{*}\to\infty$, which is as follows (Dinc et al., 2025b):
@@
These expressions differ from (Dinc et al., 2025b) only in the sense that the normalization of the loss function introduces a $(T)^{-1}$ coefficient to the gradient computation. Notably, this loss function has a kink at its global optimum. Thanks to this kink, (Dinc et al., 2025b) has defined a maximum (effective) learning rate $\alpha^{*}$, for any $\alpha>\alpha^{*}$, the model would be thrown from the global minimum to the zero gradient regime, after which the learning would come to an halt despite a non-desirable loss value of $1$. This learning rate can be computed as:
@@
S2.1.2 Ghost model with asymmetric response time
@@
As shown by previous work (Dinc et al., 2025b) (by integrating the dynamical system equations), in the limit $x^{*}\to\infty$, the network output can be written as $\hat{o}(t)=\Theta(t-t^{*})$, where $t^{*}=\frac{\pi}{2\sqrt{r}}$. This creates three distinct regimes of analytical calculation: i) $t^{*}\leq T_{\rm delay}$, ii) $T_{\rm delay}\leq t^{*}\leq T_{\rm delay}+T_{\rm resp}$, and iii) $T_{\rm delay}+T_{\rm resp}\leq t^{*}$.
@@
$\mathcal{L}(r)=\begin{cases}1\quad&\text{for}\quad r\leq\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm resp})^{2}},\\ \frac{\pi}{2T_{\rm resp}\sqrt{r}}-\frac{T_{\rm delay}}{T_{\rm resp}}\quad&\text{for}\quad\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm resp})^{2}}\leq r\leq\frac{\pi^{2}}{4T_{\rm delay}^{2}},\\ 1-\frac{\pi}{2T_{\rm delay}\sqrt{r}}\quad&\text{for}\quad r\geq\frac{\pi^{2}}{4T_{\rm delay}^{2}}.\end{cases}$ (S11)
@@
Similar to before, one can compute the critical learning rate due to the kink at the global optimum. First, noting $r_{b}=\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm resp})^{2}}$, we arrive at:
$r^{*}-r_{b}=\frac{\pi^{2}}{4T_{\rm delay}^{2}}-\frac{\pi^{2}}{4(T_{\rm delay}+T_{\rm resp})^{2}}=\frac{\pi^{2}(2T_{\rm delay}+T_{\rm resp})T_{\rm resp}}{4T_{\rm delay}^{2}(T_{\rm delay}+T_{\rm resp})^{2}}\xrightarrow{T_{\rm delay}\gg T_{\rm resp}}\frac{\pi^{2}T_{\rm resp}}{2T_{\rm delay}^{3}}.$ (S13)
@@
S2.2 Limit cycle model
@@
S2.2.2 Analytical calculations with equal delay and response times
@@
S2.2.3 Analytical calculations with long delays
@@
Appendix S3 Experimental details on empirical evaluations
S3.1 Simulating RNN models
In this section, we describe how we implemented the update rule for the leaky firing-rate RNN defined in Equation (1). Since the experiments are conducted in discrete time, we use the forward Euler method to update the firing rates:
$r(t+1)=r(t)+\alpha f(r(t),u(t))$ (S30)
Here, $r$ denotes the firing rate, $f(\cdot)$ corresponds to the update function defined in Equation (1), and $\alpha$ is the step size, computed as $\alpha=\Delta t/\tau$. The update function can be written more explicitly as:
$r(t+1)=(1-\alpha)\,r(t)+\alpha\,\tanh(Wr(t)+W^{in}u(t)+b+\epsilon)$ (S31)
With this update rule, we can compute the firing rate at the next time step. The only remaining component is computing the gradients from the predicted output $\hat{o}(t)$ to the target output $o^{*}(t)$, as described in Sections S3.2 and S3.3.
S3.2 Task details
Short-term memory is a fundamental cognitive process essential for mammalian survival. Several well-established studies have proposed theories on memory maintenance that have deepened our understanding of short-term memory, as illustrated in Fig. 1. However, while these theories focus solely on memory maintenance and response processes, they do not fully explain the underlying mechanisms. This knowledge gap raises a critical question: Which mechanisms govern short-term memory in the brain? To address this question, we investigate short-term memory through two key properties: memory maintenance and information processing. Accordingly, we conduct experiments on two well-established neuroscience tasks: the delayed activation and delayed cue-discrimination tasks. We determined these tasks because the first task exclusively examines memory maintenance, while the second incorporates both information processing and the retention of input stimuli to generate the relevant response. Furthermore, to enhance transparency and ensure the reproducibility of our experiments, we provide a detailed outline of all task properties in this section.
S3.2.1 Delayed Activation Task
The delayed activation task allows us to isolate the memory component by assuming that cue processing has already occurred. In this setup, the task consists of three distinct periods: the delay period $(T_{\rm delay})$, the reaction period $(T_{\rm resp})$, and the optional post-reaction period $(T_{\rm post})$. Since the cue period is omitted, the input remains zero throughout the trial, making the task entirely dependent on the intrinsic dynamics of the recurrent model rather than external stimuli.
The expected ground truth responses, denoted as $O^{*}=\{o_{1}^{*},o_{2}^{*},o_{3}^{*},...,o_{T}^{*}\}$, are defined over the total duration $T=T_{\rm delay}+T_{\rm resp}+T_{\rm post}$ as follows:
$o_{t}^{*}=\begin{cases}1,&\text{if }t\in T_{\rm resp}\\ 0,&\text{otherwise}\end{cases}$ (S32)
This task is inherently simpler than the delayed cue-discrimination task, as it isolates the challenge of maintaining information in memory without requiring cue processing. By setting the input $u$ to zero at all times, the model must rely solely on its internal state dynamics to correctly generate the expected response during the reaction period. The optional post-reaction period allows for additional investigation into how the network dynamics stabilize after completing the required response.
S3.2.2 Delayed Cue-Discrimination Task
In contrast to the delayed activation task, the delayed cue-discrimination task involves two input channels and two output channels. It consists of four distinct periods: the cue period $(T_{\rm in})$, delay period $(T_{\rm delay})$, reaction period $(T_{\rm resp})$, and an optional post-reaction period $(T_{\rm post})$. This task allows us to investigate both stimulus processing and memory components simultaneously.
The input cue for the delayed cue-discrimination task, denoted as
$U=\{(u_{1}^{1},u_{1}^{2}),(u_{2}^{1},u_{2}^{2}),(u_{3}^{1},u_{3}^{2}),...,(u_{T}^{1},u_{T}^{2})\}$
is defined over the total duration $T=T_{\rm in}+T_{\rm delay}+T_{\rm resp}+T_{\rm post}$ as follows:
$u_{t}=\begin{cases}(0,1)\text{ or }(1,0),&\text{if }t\in T_{\rm in}\\ (0,0),&\text{otherwise}\end{cases}$ (S33)
The expected ground truth outputs, denoted as
$O^{*}=\{(o_{1}^{*1},o_{1}^{*2}),(o_{2}^{*1},o_{2}^{*2}),(o_{3}^{*1},o_{3}^{*2}),...,(o_{T}^{*1},o_{T}^{*2})\}$
are defined as:
$o_{t}^{*}=\begin{cases}u_{T_{in}},&\text{if }t\in T_{\rm resp}\\ (0,0),&\text{otherwise}\end{cases}$ (S34)
This task requires the model to process the cue presented during the cue period, maintain it through the delay period, and then generate the appropriate response during the reaction period. The optional post-reaction period allows for additional investigation into how the network dynamics stabilize after completing the required response.
S3.3 Training details
To ensure the reproducibility of our experiments, we provide a comprehensive description of the training configurations used in this study. All recurrent neural network (RNN) architectures were implemented using the PyTorch framework. Unless stated otherwise, models were trained using the stochastic gradient descent (SGD) with the momentum value of $0.9$ and weight decay value of $10^{-7}$. Most of the experiments were conducted on computing systems equipped with an Intel i9-10900X CPU and an Apple M2 CPU. To further enhance transparency and support future research, we will publicly release our complete training pipeline on GitHub.
Training with long delay intervals, i.e., $T_{\rm delay}\gg T_{\rm resp}$, presents challenges when compared to shorter delay times due to the model’s tendency to learn to always output 00. This issue arises from an inherent imbalance in the loss function: since the model is expected to output 00 for the majority of the trial and only $1$ during the response period, a naive mean-squared error (MSE) loss would favor minimizing errors in the dominant interval, leading to trivial solutions where the model outputs 00 at all times. This imbalance is aggravated when the total trial duration $T_{\rm in}+T_{\rm delay}+T_{\rm post}$ is significantly greater than $T_{\rm resp}$.
To mitigate this issue, we trained all RNNs using a weighted MSE loss, where different time intervals were assigned corresponding weights. The weight function was defined as follows:
$w_{MSE}=\begin{cases}1-\frac{10}{T_{\text{total}}},&\text{if }t\in T_{\rm resp}\\ \frac{10}{T_{\text{total}}},&\text{otherwise}\end{cases}$ (S35)
This weighting scheme ensures that the loss function does not disproportionately favor the dominant 00 responses and appropriately emphasizes the importance of correct responses during the reaction period.
S3.4 Details on large-scale experiments
In this study, we conducted several short-term memory experiments to investigate underlying memory mechanisms. Specifically, we examined rank-2 RNNs trained on the delayed activation task, both with and without a post-reaction period (Fig. 2, 3, and S1). To extend our observations to full-rank RNNs, we increased task complexity by training models on the delayed cue-discrimination task, again with and without a post-reaction period (Fig. 4, S3, and S4). Additionally, to explore the role of the time constant $\tau$ and memory-related components, we evaluated full-rank RNNs under various configurations (Fig. S5 and S6). Further implementation and training details are provided in Sections S4 and S3.3. As a summary, the configurations and number of total runs for each experiment are presented below:
Table S1: Large-scale experiment configurations and the number of total runs.
RNN Task Learning Rate $T_{\rm delay}$ Seeds # Experiments
Rank-2 Delayed act. w/wo $T_{\rm post}$ $5\times 10^{-3}$ $30\,\mathrm{ms}$ 100 1600
Full Cue-disc. wo/ $T_{\rm delay}$ $\log[10^{-3},1]$ (20) $T_{\rm resp}=\mathrm{lin}[20,400]\,\mathrm{ms}$ (20) 20 8000
Full Cue-disc. w/ longer $\tau$ $\log[10^{-3},10^{-1}]$ (10) $\mathrm{lin}[0,800]\,\mathrm{ms}$ (21) 20 4200
Full Cue-disc. wo/ $T_{\rm post}$ $\log[10^{-3},10^{-1}]$ (15) $\mathrm{lin}[0,400]\,\mathrm{ms}$ (21) 100 31500
Full Cue-disc. wo/ $T_{\rm post}$ $\log[10^{-1.6},10^{-0.5}]$ (15) $\mathrm{lin}[0,400]\,\mathrm{ms}$ (21) 100 31500
Full Cue-disc. w/ $T_{\rm post}$ $\log[10^{-3},10^{-1}]$ (15) $\mathrm{lin}[0,400]\,\mathrm{ms}$ (21) 20 6300
S3.4.1 Detecting learned mechanisms subserving short-term memory in RNNs
We designed two metrics—Reaction Accuracy (RA) and Reaction Reliability (RR)—to evaluate the performance of our recurrent neural network (RNN) models. Reaction Accuracy quantifies the model’s ability to discriminate the input cue by identifying the class corresponding to the maximum output channel at each time step, while Reaction Reliability assesses the degree of signal distinction between the classes.
More formally, let $O\in\mathbb{R}^{N\times 2}$ represent the model’s output vector over the reaction interval, and $G\in\mathbb{R}^{N\times 1}$ denote the ground truth labels over the same interval, where $N$ is the total number of time steps.
Reaction Accuracy (RA):
This metric assesses how accurately the model produces the correct response during the response period $(T_{\rm resp})$. It is computed as follows:
$RA=1-\frac{1}{N}\sum_{t=1}^{N}\left|\arg\max_{c}(O_{t}^{c})-G_{t}\right|$ (S36)
Here, $\arg\max_{c}(O_{t}^{c})$ returns the index of the output channel with the highest activation at time step $t$, and $G_{t}$ denotes the ground truth class label at that time.
Reaction Reliability (RR):
This metric complements Reaction Accuracy by not only evaluating whether the model selects the correct response during the response period $T_{\rm resp}$, but also by measuring how confidently it does so. It is computed as follows:
$RR=\frac{1}{N}\sum_{t=1}^{N}\frac{O_{t,G_{t}}}{\sum_{c=1}^{2}O_{t}^{c}}$ (S37)
Here, $O_{t,G_{t}}$ denotes the model’s output corresponding to the ground truth class $G_{t}$ at time step $t$, and the denominator normalizes this value by the total output across both output channels at that time step.
Mechanism Discrimination Index (MDI):
The two metrics, RA and RR, allow us to determine whether the model has successfully learned the task. If a model fails to learn, we categorize its behavior as either ‘Unstable Training’ or ‘Final Model Failure’, as illustrated in Fig. S3A-B and Fig. S4A. However, these metrics alone do not clearly distinguish between limit cycles and slow-point manifolds (see Fig. S3A-B and Fig. S4C-D). Therefore, to reliably differentiate these two mechanisms, we further analyze the frequency distributions of correlation coefficient matrices computed from the first two principal components of the models’ firing rates, as shown in Fig. S3C.
Since limit cycles closely approximate slow-point manifolds during the trial period, it is challenging to distinguish them based solely on firing rates observed within that interval. To overcome this limitation, we extend the analysis window by concatenating the post-reaction period during inference, thereby revealing the underlying dynamical structure of the firing rates, as illustrated in Fig. S4C-D. Next, we apply principal component analysis (PCA) to the firing rates and extract their first two principal components (Fig. S4). We then compute the Pearson product-moment correlation coefficient matrix from these principal components.
Due to the repetitive structure of limit cycles, their correlation coefficient matrices exhibit predominantly high-frequency components. In contrast, slow-point manifolds produce correlation matrices characterized by consistently high correlations, corresponding predominantly to low-frequency components. After calculating the frequency components, we define a hyperparameter to establish a threshold for discrimination between low-frequency and high-frequency, and compute the Mechanism Discrimination Index (MDI) as follows:
$MDI=\frac{PSD_{lf}}{PSD_{total}}$ (S38)
where $PSD_{lf}$ is the power spectral density of low-frequency components, and $PSD_{total}$ is the total power spectral density. Using this approach, we reliably discriminate between limit cycles and slow-point manifolds, as demonstrated in Fig. S3C.
As a summary, we defined four categories to classify the outcomes of RNN training:
  • Final Model Failure: An RNN is considered to have failed to converge if its mean accuracy and reliability have never previously reached or exceeded $0.8$ (Fig. S4A).